When a rabbit fur is rubbed with a acetate fiber, the fur acquires a change of magnitude 2.23nC.


a. Did the rabbit fur gain or lose electrons?

b. How many electrons were transferred during the process?

c. What is the change in mass of the rabbit fur?

d. What is the change in mass of the acetate fiber?



1
Expert's answer
2022-01-18T09:35:59-0500

a. Fur loses electrons.

b. N=Q/e=1.39·1010.

c. m=meN=1.27·10-20 kg.

d. Here, the change in mass is the same as in (c) above.